# Real-Time Dynamics in Two Dimensions with Tensor Network States via Time-Dependent Variational Monte Carlo

### v3

@@ Abstract
Reliably simulating two-dimensional many-body quantum dynamics with projected entangled pair states (PEPS) has long been a difficult challenge. In this work, we overcome this barrier for low-energy quantum dynamics by developing a stable and efficient time-dependent variational Monte Carlo (tVMC) framework for PEPS. By analytically removing all gauge redundancies of the PEPS manifold and exploiting tensor locality, we obtain a numerically well-conditioned stochastic reconfiguration equation amenable to robust solution using the efficient Cholesky decomposition, enabling long-time evolution in previously inaccessible regimes. We explain how the difficulties associated with the traditional PEPS algorithms for dynamics are naturally resolved within tVMC. We demonstrate the power and generality of the method through four representative real-time problems in two dimensions: (I) chiral edge propagation in a free-fermion Chern insulator; (II) fractionalized charge transport in a fractional Chern insulator; (III) vison confinement dynamics in the Higgs phase of a $\mathbb{Z}_{2}$ lattice gauge theory coupled to a Higgs field; and (IV) superfluidity and critical velocity in interacting bosons. All simulations are performed on $12\times 12$ or $13\times 13$ lattices with evolution times $T=10$ to $12$ using modest computational resources — 1 to 5 days on a single GPU card. Where exact benchmarks exist (case I), PEPS–tVMC matches free-fermion dynamics with high accuracy up to $T=12$. These results demonstrate PEPS–tVMC as a practical and versatile tool for real-time quantum dynamics in two dimensions. The method significantly extends the reach of classical tensor-network simulations for studying elementary excitations in quantum many-body systems in real-time and provides a valuable computational counterpart to emerging quantum simulators.
@@ I Introduction
Understanding real-time dynamics of interacting quantum many-body systems in two dimensions (2D) is a fundamental challenge that spans condensed-matter physics, particle physics, and quantum simulation. While powerful classical methods exist for ground states—including tensor networks [1, 2] and quantum Monte Carlo [3, 4], real-time evolution in 2D remains notoriously difficult. Exact diagonalization is limited to small clusters; quantum Monte Carlo suffers from a dynamical sign problem. Recently, the combination of time-dependent variational Monte Carlo (tVMC) [5, 6, 7], a continuous-time generalization of the stochastic reconfiguration method [8, 9, 10], with neural quantum states (NQS) [11] has emerged as a promising alternative [12, 13, 14, 15, 16, 17, 18]. NQS provide remarkable variational flexibility, while tVMC offers a principled formulation of real-time evolution. At present, extending NQS–tVMC to large 2D systems remains an active area of research, where one must balance expressive parametrizations with efficient and stable optimization [19].
@@
Projected entangled pair states (PEPS) [20] provide a complementary variational family whose locality and multi-linearity make them particularly appealing for tVMC. Within VMC, only a single layer of PEPS tensors needs to be contracted, making the method substantially more efficient than conventional PEPS update schemes [21, 22, 23]. Furthermore, VMC allows parallel accumulation of samples, making the approach well suited for modern GPU architectures. In fact, PEPS-VMC has recently achieved considerable success for ground states, giving competitive results for the ground states of frustrated spins [24, 25, 26], fermion Hubbard model [27], and lattice gauge theory [28], owing to a foundational breakthrough in the sampling method of PEPS [29].
@@
Although PEPS obey an area law for entanglement entropy [30] and cannot represent volume-law states produced by fully thermalizing dynamics, they are well suited for local quenches in which the injected energy does not scale with system size. In this regime, the energy density remains zero (in the thermodynamic limit) and we expect the dynamics to preserve area-law entanglement [31, 32]. Thus PEPS naturally targets low-energy local-quench dynamics — precisely the regime relevant for elementary excitations and real-time response functions, and directly related to experiments including scanning tunneling microscopy [33], angle resolved photoemission spectroscopy [34], particle scattering processes, and more. A controlled method capable of simulating such dynamics provides a valuable tool for studying 2D quantum systems.
@@
While dynamical simulations with matrix product states (MPS) are enormously successful in 1D [35, 36, 37, 38], the situation in 2D is far more challenging. In contrast to the extensive 1D literature, only a small number of works—represented by Refs. [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49], have reported real-time simulations using PEPS or iPEPS, illustrating the general difficulty of the problem. Traditional approaches to dynamical PEPS have relied on full-update-like methods based on double-layer contraction [21, 22, 23], which face several challenges: (a) high cost of double-layer contraction, (b) difficulty in fixing the PEPS gauge freedom, and (c) loss of positivity in the quantum metric approximately obtained from the contraction of the double-layer environment. In Sec. III.4, we explain that, within PEPS–tVMC, the last two issues are naturally resolved while the first one is significantly alleviated, at the expense of manageable statistical noise.
@@
Despite the long development of PEPS [1], their combination with tVMC for real-time dynamics has received virtually no attention, aside from our recent exploratory attempt in a simple setting [28]. Here, we develop PEPS–tVMC systematically, analyze its structure, and demonstrate its ability to simulate large-scale 2D dynamics. Our approach builds on stochastic reconfiguration (SR) [8], identifies and removes the gauge redundancy of PEPS parameters, and exploits tensor locality to simplify and precondition the SR equations. After gauge removal, the SR equation exhibits remarkable numerical stability: the SR residual remains as low as $10^{-9}$ to $10^{-25}$ in the examples studied. The central message of this work is that by removing PEPS gauge freedom and exploiting locality, the tVMC equations become remarkably stable to solve, and that, as a result, PEPS—long viewed as difficult for real-time evolution—can in fact provide a powerful and controllable variational framework for 2D dynamics. This resolves long-standing challenges that have hindered dynamical simulations with PEPS since their introduction [20].
We illustrate the versatility of PEPS–tVMC through four physically motivated examples. First, we simulate the propagation of chiral edge modes in a Chern insulator and visualize their unidirectional boundary motion. Second, we track quantized charge transport in a fractional quantum Hall state, revealing the real-time flow of fractional charge. Third, we study vison dynamics in the Higgs phase of a $\mathbb{Z}_{2}$ lattice gauge theory, observing vison confinement in real time. Finally, we investigate superflow in interacting bosons around a local obstacle and observe signatures consistent with a critical velocity. All simulations are performed on $12\times 12$ or $13\times 13$ lattices with evolution times $T=10$ to $12$, using a single GPU card. These examples illustrate that the method is not specialized to a single model class.
@@ II Time-Dependent Variational Monte Carlo
Consider a many-body lattice quantum system with computational basis $\bm{\mathbf{s}}=(s_{1},\dots,s_{N})$. To perform a variational simulation, we adopt a holomorphic parametrization of the unnormalized quantum state $\Psi_{\bm{\mathbf{\theta}}}(\bm{\mathbf{s}})$ with $N_{p}$ complex variational parameters $\bm{\mathbf{\theta}}=\{\theta_{\alpha}:\alpha=1,2,\dots,N_{p}\}$. At time $t$, the exact evolution under a Hamiltonian $H$ for a short interval $dt$ induces an evolution from $\bm{\mathbf{\theta}}(t)$ to $\bm{\mathbf{\theta}}(t+dt)$ that minimizes the Fubini–Study distance [50] between $\Psi_{\bm{\mathbf{\theta}}(t+dt)}$ and $e^{-iHdt}\Psi_{\bm{\mathbf{\theta}}(t)}$. Taking $dt$ to 0, this yields the SR equation for tVMC [5]:
@@
where the SR matrix $S_{\alpha\beta}$ and the energy gradient $g_{\alpha}$ are given by
@@
Here $\cdot$] denotes a statistical average over configurations drawn from $P_{\bm{\mathbf{\theta}}}(\bm{\mathbf{s}})=|\Psi_{\bm{\mathbf{\theta}}}(\bm{\mathbf{s}})|^{2}/\sum_{\bm{\mathbf{s}}}|\Psi_{\bm{\mathbf{\theta}}}(\bm{\mathbf{s}})|^{2}$, $H_{\text{loc}}(\bm{\mathbf{s}})=\braket{\bm{\mathbf{s}}|H|\Psi}/\braket{\bm{\mathbf{s}}|\Psi}$ is the local energy, and
@@
the SR matrix and the energy gradient can be approximated as
@@
where the approximation reflects finite-sample statistical fluctuations. The centering in Eq. 5 removes the component parallel to the wavefunction, ensuring projective invariance of the SR metric.
The tVMC algorithm therefore consists of: (1) sampling configurations $\bm{\mathbf{s}}\sim P_{\bm{\mathbf{\theta}}}$; (2) constructing $S$ and $\bm{\mathbf{g}}$; (3) solving Eq. (1) for $\dot{\bm{\mathbf{\theta}}}$; and (4) integrating the parameter flow on the variational manifold via a Runge–Kutta method [51].
A major challenge in tVMC is that the SR matrix $S$ is often severely ill-conditioned, making Eq. (1) difficult to solve stably. For ground-state searches, where the precise time trajectory is irrelevant, one may add a diagonal regulator $S\to S+\epsilon I$ with $\epsilon\sim 10^{-3}$. However, in real-time dynamics a large $\epsilon$ introduces an unphysical bias, whereas too small a value leaves the system unstable. In NQS-tVMC, several sophisticated regularization schemes have been proposed [14, 17], but none fully resolves the problem [19].
In PEPS, the situation is complicated by gauge redundancy, which introduces many null vectors into $O$ and hence into $S$. The crucial feature, however, is that these null directions are analytically identifiable. As we will show in Sec. III, they can be removed before solving the SR equation, providing an exact preconditioning. After this gauge-removal step, the SR equation becomes remarkably well conditioned, and in all examples studied we find it can be solved with exceptional numerical accuracy. This stability is a key factor of the success of PEPS–tVMC.
@@ III Stochastic reconfiguration with PEPS
In this section we derive the SR equations for PEPS, analyze the structure of their null space, and describe how gauge redundancy can be eliminated efficiently in closed form, in both the SR and its reduced variant minSR [52] form.
Figure 1: PEPS structure and gauge freedom. (a) A $3\times 3$ PEPS network with local tensors AA. (b) A single tensor AA with physical index $p$ and virtual indices $(l,r,d,u)$. (c) Local gauge transformation on a virtual bond: inserting an invertible matrix $X$ and its inverse $X^{-1}$ between two adjacent tensors leaves the global PEPS wavefunction invariant. (d) Wavefunction amplitude $\braket{\bm{\mathbf{s}}|\Psi}$, computed via the contraction of the single layer network of sliced PEPS tensors: $a[\bm{\mathbf{x}}]_{lrdu}=A[\bm{\mathbf{x}}]^{p=\bm{\mathbf{s}}(\bm{\mathbf{x}})}_{lrdu}$.
A PEPS defines a variational wavefunction $\Psi_{\bm{\mathbf{\theta}}}(\bm{\mathbf{s}})$ parametrized by the tensor elements $A[\bm{\mathbf{x}}]^{p}_{lrdu}$ at each lattice site $\bm{\mathbf{x}}$, as illustrated in Fig. 1(a,b). Each index tuple $(\bm{\mathbf{x}},p,l,r,d,u)$ corresponds to one variational parameter, replacing the abstract index $\alpha$ introduced earlier. The physical index $p$ labels the local Hilbert-space configuration (e.g., $p=0,1$ for vacuum or occupation), while the virtual indices $(l,r,d,u)$ take values from $1$ to $D$ (except on boundaries). The PEPS in this paper all have open boundary condition.
@@
Hence every gauge vector also lies in the null space of $S=\mathcal{O}^{\dagger}\mathcal{O}$, and must be removed before solving the SR equation.
@@
III.1 Gauge-fixing within SR
We now outline the procedure for eliminating null vectors within the SR formulation. First, we construct a projection matrix onto the subspace orthogonal to the constraint vectors $\bm{\mathbf{w}}$, thereby reducing the number of independent gauge vectors to
@@ III.1 Gauge-fixing within SR
The SR equation is then solved for $y$ in the reduced space:
@@ III.2 Gauge-fixing within minSR
When the number of samples is less than the column rank of $\mathcal{O}$, i.e. $N_{s}<N_{p}-N_{\rm gv}-2$, the SR equation can be equivalently solved in the reduced minSR form [52]:
@@
Finally, we comment on the choice between SR and minSR. As the system size increases, $N_{p}$ eventually overtakes $N_{s}$ and one should use minSR for efficiency. However, it is always advised to study a problem first at small size and ideally benchmark against exact methods, in which case typically $N_{s}>N_{p}-N_{\text{gv}}-2$ and it is only correct to use SR.

### v4

@@ Abstract
Reliably simulating two-dimensional many-body quantum dynamics with projected entangled pair states (PEPS) has long been a difficult challenge. In this work, we overcome this barrier for low-energy quantum dynamics by developing a stable and efficient time-dependent variational Monte Carlo (tVMC) framework for PEPS. By analytically removing all gauge redundancies of the PEPS manifold and exploiting tensor locality, we obtain a numerically well-conditioned quantum geometric tensor, yielding robust solutions in time-dependent variational Monte Carlo using the efficient Cholesky decomposition. This enables long-time evolution in previously inaccessible regimes. We explain how the difficulties associated with the traditional PEPS algorithms for dynamics are naturally resolved within tVMC. We demonstrate the power and generality of the method through four representative real-time problems in two dimensions: (I) chiral edge propagation in a free-fermion Chern insulator; (II) fractionalized charge transport in a fractional Chern insulator; (III) vison confinement dynamics in the Higgs phase of a $\mathbb{Z}_{2}$ lattice gauge theory coupled to a Higgs field; and (IV) superfluidity and critical velocity in interacting bosons. All simulations are performed on $12\times 12$ or $13\times 13$ lattices with evolution times $T=10$ to $12$ using modest computational resources — 1 to 5 days on a single GPU card. Where exact benchmarks exist (case I), PEPS–tVMC matches free-fermion dynamics with high accuracy up to $T=12$. These results demonstrate PEPS–tVMC as a practical and versatile tool for low-energy real-time quantum dynamics in two dimensions. The method significantly extends the reach of classical tensor-network simulations for studying elementary excitations in quantum many-body systems in real-time and provides a valuable computational counterpart to emerging quantum simulators.
@@ I Introduction
Understanding real-time dynamics of interacting quantum many-body systems in two dimensions (2D) is a fundamental challenge that spans condensed-matter physics, particle physics, and quantum simulation. While powerful classical methods exist for ground states—including tensor networks [1, 2] and quantum Monte Carlo [3, 4], real-time evolution in 2D remains notoriously difficult. Exact diagonalization is limited to small clusters; quantum Monte Carlo suffers from a dynamical sign problem. Recently, the combination of time-dependent variational Monte Carlo (tVMC) [5, 6, 7, 8], a continuous- and real-time generalization of the stochastic reconfiguration method [9, 10, 11], with neural quantum states (NQS) [12] has emerged as a promising alternative [13, 14, 15, 16]. Here, we refer to tVMC as a Monte Carlo formulation of the time-dependent variational principle (TDVP) [5, 6]. NQS provides remarkable variational flexibility [17, 18, 19], while tVMC offers a principled formulation of real-time evolution. The combination has also motivated alternative approaches to dynamics within the VMC framework [14, 20, 16, 21]. At present, extending NQS–tVMC to large 2D systems remains an active area of research, where one must balance expressive parametrizations with efficient and stable optimization [22].
@@
Projected entangled pair states (PEPS) [23] provide a complementary variational family whose locality and multi-linearity make them particularly appealing for tVMC. Within VMC, only a single layer of PEPS tensors needs to be contracted, making the method substantially more efficient than conventional PEPS update schemes [24, 25, 26]. Furthermore, VMC allows parallel accumulation of samples, making the approach well suited for modern GPU architectures. In fact, PEPS-VMC has recently achieved considerable success for ground states, giving competitive results for the ground states of frustrated spins [27, 28, 29], the Fermi-Hubbard model [30], and lattice gauge theory [31], owing to a foundational breakthrough in the sampling method of PEPS [32].
@@
Although PEPS obey an area law for entanglement entropy [33] and cannot represent volume-law states produced by fully thermalizing dynamics, they are well suited for local quenches in which the injected energy does not scale with system size. In this latter regime, the energy density remains zero (in the thermodynamic limit) and we expect the dynamics to preserve area-law entanglement [34, 35]. Thus, PEPS naturally targets low-energy local-quench dynamics — precisely the regime relevant for elementary excitations and real-time response functions, and directly related to experiments including scanning tunneling microscopy [36], angle-resolved photoemission spectroscopy [37], particle scattering processes, and more. A controlled method capable of simulating such dynamics provides a valuable tool for studying 2D quantum systems.
@@
While dynamical simulations with matrix product states (MPS) are enormously successful in 1D [38, 39, 40, 41], the situation with PEPS in 2D is far more challenging. In contrast to the extensive 1D literature, only a small number of works—represented by Refs. [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52], have reported real-time simulations using PEPS or iPEPS, illustrating the general difficulty of the problem. Traditional approaches to dynamical PEPS have relied on full-update-like methods based on double-layer contraction [24, 25, 26], which face several challenges: (a) high cost of double-layer contraction, (b) difficulty in fixing the PEPS gauge freedom, and (c) loss of positivity in the quantum metric approximately obtained from the contraction of the double-layer environment. In Sec. III.4, we explain that, within PEPS–tVMC, the last two issues are naturally resolved while the first one is significantly alleviated, at the expense of manageable statistical noise.
@@
Despite the long development of PEPS [1], their combination with tVMC for real-time dynamics has received virtually no attention, aside from our recent exploratory attempt in a simple setting [31]. Here, we develop PEPS–tVMC systematically, analyze its structure, and demonstrate its ability to simulate large-scale 2D dynamics. Our approach builds on time-dependent Variational Monte Carlo  [5, 6], and identifies and removes the gauge redundancy of PEPS parameters and exploits tensor locality to simplify and precondition the TDVP equations. After gauge removal, the TDVP equation exhibits remarkable numerical stability: the TDVP residual remains as low as $10^{-9}$ to $10^{-25}$ in the examples studied. As a result, PEPS—long viewed as difficult for real-time evolution—can in fact provide a powerful and controllable variational framework for 2D low-energy dynamics. This resolves long-standing challenges that have hindered dynamical simulations with PEPS since their introduction [23]. We illustrate the versatility of PEPS–tVMC through four physically motivated examples, which illustrate that the method is not specialized to a single model class.
@@ II Time-Dependent Variational Monte Carlo
Consider a many-body lattice quantum system with computational basis $\bm{\mathbf{s}}=(s_{1},\dots,s_{N})$ with local Hilbert space dimension $d$. To perform a variational simulation, we adopt a holomorphic parametrization of the unnormalized quantum state $\Psi_{\bm{\mathbf{\theta}}}(\bm{\mathbf{s}})$ with $N_{p}$ complex variational parameters $\bm{\mathbf{\theta}}=\{\theta_{\alpha}:\alpha=1,2,\dots,N_{p}\}$. At time $t$, the exact evolution under a Hamiltonian $H$ for a short interval $dt$ induces an evolution from $\bm{\mathbf{\theta}}(t)$ to $\bm{\mathbf{\theta}}(t+dt)$ that minimizes the Fubini–Study distance [53] between $\Psi_{\bm{\mathbf{\theta}}(t+dt)}$ and $e^{-iHdt}\Psi_{\bm{\mathbf{\theta}}(t)}$. For infinitesimal $dt$, this yields the TDVP equations [5, 38]:
@@
where the gram matrix or quantum geometric tensor (QGT) $S_{\alpha\beta}$ and the energy gradient $g_{\alpha}$ are given by (up to normalization factors)
$S_{\alpha\beta}$ $=\braket{\partial_{\alpha}\Psi|\partial_{\beta}\Psi}-\braket{\partial_{\alpha}\Psi|\Psi}\braket{\Psi|\partial_{\beta}\Psi},$ (2)
$g_{\alpha}$ $=\braket{\partial_{\alpha}\Psi|H|\Psi}-\braket{\partial_{\alpha}\Psi|\Psi}\braket{\Psi|H|\Psi}.$ (3)
These quantities can be written in terms of expectation values over the Born distribution $P_{\bm{\mathbf{\theta}}}(\bm{\mathbf{s}})=|\Psi_{\bm{\mathbf{\theta}}}(\bm{\mathbf{s}})|^{2}/\sum_{\bm{\mathbf{s}}}|\Psi_{\bm{\mathbf{\theta}}}(\bm{\mathbf{s}})|^{2}$
@@
Here $\cdot$] denotes a statistical average over $P_{\bm{\mathbf{\theta}}}(\bm{\mathbf{s}})$, $H_{\text{loc}}(\bm{\mathbf{s}})=\braket{\bm{\mathbf{s}}|H|\Psi}/\braket{\bm{\mathbf{s}}|\Psi}$ is the local energy, and
@@
the QGT and the energy gradient can be approximated as
@@
where the approximation reflects finite-sample statistical fluctuations. The centering in Eq. (7) removes the component parallel to the wavefunction, ensuring projective invariance of the quantum metric.
The tVMC algorithm therefore consists of: (1) sampling configurations $\bm{\mathbf{s}}\sim P_{\bm{\mathbf{\theta}}}$; (2) constructing $S$ and $\bm{\mathbf{g}}$; (3) solving Eq. (1) for $\dot{\bm{\mathbf{\theta}}}$; and (4) integrating the parameter flow on the variational manifold via a Runge–Kutta integrator [54].
A major challenge in tVMC is that the QGT $S$ is often severely ill-conditioned, making the linear problem in Eq. (1) difficult to solve stably. For ground-state searches, where the precise time trajectory is irrelevant, one may add Tikhonov regularization $S\to S+\epsilon I$ with typically $\epsilon\sim 10^{-3}$. However, in real-time dynamics a large $\epsilon$ introduces an unphysical bias, whereas too small a value leaves the system unstable. In NQS-tVMC, several sophisticated regularization schemes have been proposed [13, 15], but none fully resolve the problem [22].
In PEPS, the main source of instability can be identified by the gauge redundancy, which introduces many null vectors into $O$ and hence into $S$. The crucial feature, however, is that these null directions are analytically identifiable. As we will show in Sec. III, they can be removed before solving the TDVP equation, providing an exact preconditioning. After this gauge-removal step, the TDVP equation becomes remarkably well conditioned, and in all examples of local quench dynamics studied we find it can be solved with exceptional numerical accuracy. This stability is a key factor in the success of PEPS–tVMC.
@@ III Stochastic reconfiguration with PEPS
In this section, we derive the TDVP equations for PEPS, analyze the structure of their null space, and describe how gauge redundancy can be eliminated efficiently in closed form, in both the TDVP and its reduced variant based on the push-through identity (referred to as minSR) [55].
Figure 1: PEPS structure and gauge freedom. (a) A $3\times 3$ PEPS network with local tensors AA. (b) A single tensor AA with physical index $p$ and virtual indices $(l,r,d,u)$. (c) Local gauge transformation on a virtual bond: inserting an invertible matrix $X$ and its inverse $X^{-1}$ between two adjacent tensors leaves the global PEPS wavefunction invariant. (d) Wavefunction amplitude $\braket{\bm{\mathbf{s}}|\Psi}$, computed via the contraction of the single layer network of sliced PEPS tensors: $a[\bm{\mathbf{x}}]_{lrdu}=A[\bm{\mathbf{x}}]^{p=\bm{\mathbf{s}}[\bm{\mathbf{x}}]}_{lrdu}$.
A PEPS defines a variational wavefunction $\Psi_{\bm{\mathbf{\theta}}}(\bm{\mathbf{s}})$ parametrized by the tensor elements $A[\bm{\mathbf{x}}]^{p}_{lrdu}$ at each lattice site $\bm{\mathbf{x}}$, as illustrated in Fig. 1(a,b). Each index tuple $(\bm{\mathbf{x}},p,l,r,d,u)$ corresponds to one variational parameter, replacing the abstract index $\alpha$ introduced earlier. The physical index $p$ labels the local Hilbert-space configuration (e.g., $p=0,1$ for vacuum or occupation), while the virtual indices $(l,r,d,u)$ take values from $1$ to $D$ (except on boundaries). The PEPS in this paper all have open boundary conditions.
@@
Hence every gauge vector also lies in the null space of $S=\mathcal{O}^{\dagger}\mathcal{O}$, and must be removed before solving the TDVP equation.
@@
III.1 Gauge-fixing within TDVP
We now outline the procedure for eliminating null vectors within the TDVP formulation. First, we construct a projection matrix onto the subspace orthogonal to the constraint vectors $\bm{\mathbf{w}}$, thereby reducing the number of independent gauge vectors to
@@ III.1 Gauge-fixing within TDVP
The TDVP equation is then solved for $y$ in the reduced space:
@@ III.2 Gauge-fixing within minSR
When the number of samples is less than the column rank of $\mathcal{O}$, i.e. $N_{s}<N_{p}-N_{\rm gv}-2$, the TDVP equation can be equivalently solved in the dual minSR form obtained using the push-through identity [55]:
@@
Finally, we comment on the choice between TDVP and minSR form. As the system size increases, $N_{p}$ eventually overtakes $N_{s}$ and one should use minSR for efficiency. However, it is always advised to study a problem first at small size and ideally benchmark against exact methods, in which case typically $N_{s}>N_{p}-N_{\text{gv}}-2$ and it is only correct to use standard TDVP.